OpenStudy (anonymous):

The chess club has 18 students. How many different teams of 4 can be formed for competition? answers below

OpenStudy (anonymous):

3,060 18,360 73,440

Parth (parthkohli):

18c4

OpenStudy (anonymous):

?

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {18! \over (18 - 4)! \times 4!} }\) \(\Large \color{Black}{\Rightarrow {18! \over14! \times 4!} }\) \(\Large \color{Black}{\Rightarrow {18 \times 17 \times 16 \times 15 \over 4 \times 3 \times 2} }\)

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {73440 \over 24} }\) \(\Large \color{Black}{\Rightarrow 3060 }\)

Parth (parthkohli):

First option ^_^

OpenStudy (anonymous):

ok thanks